# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ Abstract
Recent literature has effectively leveraged diffusion models trained on continuous variables as priors for solving inverse problems. Notably, discrete diffusion models with discrete latent codes have shown strong performance, particularly in modalities suited for discrete compressed representations, such as image and motion generation. However, their discrete and non-differentiable nature has limited their application to inverse problems formulated in continuous spaces. This paper presents a novel method for addressing linear inverse problems by leveraging image-generation models based on discrete diffusion as priors. We overcome these limitations by approximating the true posterior distribution with a variational distribution constructed from categorical distributions and continuous relaxation techniques. Furthermore, we employ a star-shaped noise process to mitigate the drawbacks of traditional discrete diffusion models with absorbing states, demonstrating that our method performs comparably to continuous diffusion techniques. To the best of our knowledge, this is the first approach to use discrete diffusion model-based priors for solving image inverse problems.
@@ 1 Introduction
Diffusion models have gained significant attention as deep generative models, achieving remarkable success in image (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021; Dhariwal & Nichol, 2021; Esser et al., 2024), audio (Liu et al., 2023; Chen et al., 2024a), and video generation (Ho et al., 2022b; a). These models operate by iteratively corrupting data then learning to reverse this corruption process, ultimately generating high-quality samples from noise. In parallel with continuous diffusion models, discrete diffusion models have emerged as a compelling alternative. These models have gained traction by demonstrating notable results not only in image (Gu et al., 2022), audio (Yang et al., 2023), and text generation (Austin et al., 2021; Lou et al., 2023a) but also in more specialized areas such as motion data (Lou et al., 2023b; Pinyoanuntapong et al., 2024), protein synthesis (Gruver et al., 2024), and graph generation (Vignac et al., 2023).
@@
Building on these advancements, researchers have made significant progress in expanding the application of diffusion models. They have explored using diffusion models, trained either directly on pixel space or on latent representations derived from variational autoencoders (VAEs), to address inverse problems (Kawar et al., 2022; Chung et al., 2023; Wang et al., 2023) and carry out various conditional-generation tasks (Yu et al., 2023; Bansal et al., 2024; He et al., 2024) without the need for additional training. These efforts aim to use the powerful generative capabilities of diffusion models to tackle intricate problems and generate conditional outputs, all while preserving the models’ original trained parameters.
@@
This line of work has been primarily restricted to diffusion models trained in continuous spaces, and methods using trained discrete diffusion models as priors remain limited Gruver et al. (2024); Chen et al. (2024b); Li et al. (2024). The inherent nature of the generation process in discrete diffusion models involves non-differentiable operations, posing a challenge for their application to inverse problems formulated in continuous spaces. Therefore, controlling discrete diffusion models often necessitates an additional trained network (Gruver et al., 2024; Nisonoff et al., 2024; Klarner et al., 2024; Vignac et al., 2023). Training-free methods have been confined to relatively low-dimensional data (Chen et al., 2024b) or to specific tasks such as image inpainting (Gu et al., 2022).
This limitation constrains the utilization of the powerful generative capabilities in domains such as motion data, where generative models with discrete latent spaces have demonstrated remarkable success. This constraint motivates the exploration of discrete diffusion models as priors, given their potential advantages in representing complex data distributions and generating high-fidelity samples. To the best of our knowledge, in the context of inverse problems for motion data and image data, there exists no prior work that uses discrete diffusion models as priors.
Controlling the generation of diffusion models without additional training generally involves manipulating the generation trajectory using gradients in continuous space. The primary objective here is to generate samples from the prior model that have high likelihood with respect to the measurement equation of the inverse problem or the guidance target. A specific approach to achieve this involves adjusting samples during the diffusion model’s generation process by using gradients of a loss function that computes the likelihood. This methodology has been demonstrated in the works of Chung et al. (2023) and Yu et al. (2023).
@@
Figure 1: Illustration of G2D2. At each time step $t$, variational categorical distribution $\tilde{p}_{\bm{\alpha}}$ is optimized with respect to sum of prior loss and likelihood loss, followed by sampling $\mathbf{z}_{t-1}$. Both terms are continuously differentiable, enabling continuous optimization.
@@
We propose Gradient-guided Discrete Diffusion (G2D2), an inverse problem solving method that overcomes the aforementioned limitations while using a discrete diffusion model as a prior. Our focus is on solving image inverse problems using a generative model based on a discrete diffusion model specifically designed for images with discrete latent variables such as those found in vector-quantized (VQ)-VAE models. G2D2 overcomes the limitations of previous methods by using a continuous relaxation technique to optimize the parameters of a variational distribution, effectively bridging the gap between continuous and discrete domains.
@@
Discrete diffusion models for image data often use a variant called ”mask-absorbing” due to its efficiency. However, this model has a significant drawback in inverse-problem solving. While a substantial portion of the image structure is determined in the initial stages of generation in discrete diffusion models (i.e., when only a few tokens are determined), the mask-absorbing type does not allow transitions from an unmasked state to either a masked state or another unmasked state. Our experiments show that this restriction imposes a significant limitation on performance.
To address this issue, we use the star-shaped noise process previously proposed in the context of continuous diffusion models (Okhotin et al., 2024; Zhang et al., 2024). This process removes the dependency between consecutive sampling steps, thus expanding the solution space that can be explored. It therefore enables the correction of errors introduced in the early phases of sampling.
@@
Originally proposed to enhance the performance of diffusion models (Okhotin et al., 2024), it was later introduced as a decoupled noise annealing process in the context of inverse problems using continuous diffusion models, demonstrating its effectiveness (Zhang et al., 2024). In this study, we not only demonstrate that this process can be effectively applied to discrete diffusion models, but also find that it uniquely addresses potential issues inherent in mask-absorbing-type discrete diffusion processes.
@@
We conduct an experimental investigation to evaluate the performance of G2D2 by comparing it to current methods using standard benchmark datasets. We consider methods that use both pixel-domain and latent diffusion models. We also explore the application of a discrete prior-based motion-data-generation model to solve an inverse problem, specifically path-conditioned generation, without requiring further training. The results of our study indicate that G2D2 shows promise in tackling various inverse problems by leveraging pre-trained discrete diffusion models.
@@ 2.1 Discrete diffusion models for image generation
We first provide a brief overview of VQ-Diffusion (Gu et al., 2022; Tang et al., 2022), an image-generation model based on discrete diffusion processes. VQ-Diffusion generates images in a two-step process. It first produces discrete latent representations $\mathbf{z}_{0}$ using a discrete diffusion model trained on representations obtained from a pre-trained VQ-VAE model (Van Den Oord et al., 2017). It then transforms these representations into the continuous image space using a decoder. Each element of $\mathbf{z}_{0}\in\{1,\ldots,K\}^{d_{\mathbf{z}}}$ corresponds to one of the embedding vectors from the codebook, denoted as $\mathbf{B}:=\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K}\},\mathbf{b}_{k}\in\mathbb{R}^{d_{\mathbf{b}}}$. During decoding, a variable $\mathbf{Z}\in\mathbf{B}^{d_{\mathbf{z}}}$ is constructed through codebook assignment, where $\left(\mathbf{Z}\right)_{i}=\mathbf{b}_{z_{0,i}}$ and $z_{0,i}$ denotes the $i$-th element of $\mathbf{z}_{0}$. This variable is then fed into a continuous decoder $D:\mathbb{R}^{d_{\mathbf{b}}\times d_{\mathbf{z}}}\to\mathbb{R}^{d_{\mathbf{x}_{0}}}$ to obtain the final image: $\mathbf{x}_{0}=D(\mathbf{Z})$.
@@
where $\bm{v}(z_{t,i})$ denotes a one-hot encoded vector representing the token at time step $t$, and $Q_{t}$ represents the transition matrix, which determines the probabilities of transitions between tokens. VQ-Diffusion uses a mask-absorbing-type forward process, which introduces a special masked token denoted as [MASK] in addition to the $K$ states from the VQ-VAE. The transition matrix is defined as
@@
where the transition probabilities are determined by three parameters: $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$. Specifically, $\alpha_{t}$ represents the probability of a token remaining unchanged, $\beta_{t}$ denotes the probability of transitioning to a different unmasked token, and $\gamma_{t}$ indicates the probability of the token being replaced with the [MASK] token. The probability $\beta_{t}$ between unmasked tokens is generally set to a very small value.
@@
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})$ represents the posterior distribution determined by the forward process, and $\tilde{p}_{\theta}$ denotes the denoising network that predicts the denoised token distribution at $t$. The output of $\tilde{p}_{\theta}$ is generally modeled as independent categorical distributions for each dimension in $\mathbf{z}_{0}$. In text-to-image models such as VQ-Diffusion, $\tilde{p}_{\theta}$ is trained with text conditioning. While the true distribution $q(\mathbf{z}_{0})$ is not dimensionally independent, the whole Markov reverse process in (3) produces a distribution over categorical variables with correlations across dimensions. We distinguish between the clean distribution $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ estimated using the model and clean distribution $p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ obtained through multiple reverse diffusion steps.
@@ 2.2 Linear-inverse-problem settings
Inverse problems involve estimating unknown data from measurement data. We specifically focus on linear inverse problems in the image domain. The relationship between the measurement image $\mathbf{y}\in\mathbb{R}^{d_{\mathbf{y}}}$ and unknown ground-truth image $\mathbf{x}_{0}\in\mathbb{R}^{d_{\mathbf{x}_{0}}}$ can be represented as
@@
where $\mathbf{A}\in\mathbb{R}^{d_{\mathbf{y}}\times d_{\mathbf{\mathbf{x}_{0}}}}$ is referred to as the forward linear operator, which describes the process by which the measurement data $\mathbf{y}$ are obtained from data $\mathbf{x}_{0}$. We assume this operator is known. The term $\bm{\eta}$ represents measurement noise, which we assume follows an isotropic Gaussian distribution with a known variance $\sigma_{\bm{\eta}}^{2}$. Consequently, the likelihood function $q(\mathbf{y}|\mathbf{x}_{0})$ can be described as $\mathcal{N}(\mathbf{y};\mathbf{A}\mathbf{x}_{0},\sigma_{\bm{\eta}}^{2}\mathbf{I})$.
@@
One of the primary challenges in inverse problems is their ill-posed nature. This means that for any given measurement $\mathbf{y}$, multiple candidate solutions may exist. To address this issue and determine $\mathbf{x}_{0}$, a common approach is to assume a prior distribution for $\mathbf{x}_{0}$, such as a Laplace distribution. Diffusion models have been utilized as more powerful and expressive priors, offering enhanced capabilities in solving these inverse problems (Kawar et al., 2022; Chung et al., 2023; Wang et al., 2023; Rout et al., 2023). These methods are able to produce images that not only fit the measurement data but also exhibit high likelihood for the prior model. Given a prior $q(\mathbf{x}_{0})$, the objective in the inverse problem is to sample from the posterior distribution $q(\mathbf{x}_{0}|\mathbf{y})$, which, according to Bayes’ theorem, is proportional to $q(\mathbf{y}|\mathbf{x}_{0})q(\mathbf{x}_{0})$.
@@
These methods can be categorized based on how they incorporate the information from the measurement data $\mathbf{y}$ into the generation trajectory of diffusion models. Methods such as denoising diffusion restoration models (DDRM) (Kawar et al., 2022) and denoising diffusion null-space models (DDNM) (Wang et al., 2023) leverage the assumption of linear operators, using singular value decomposition of the forward process to control the generative process. In contrast, methods such as diffusion posterior sampling (DPS) (Chung et al., 2023) and posterior sampling with latent diffusion (PSLD) (Rout et al., 2023) operate by propagating the gradient of a loss term through the generative process. This loss term is designed to maximize the measurement likelihood, specifically by minimizing the term $\|\mathbf{y}-\mathbf{A}\mathbf{x}_{0}\|_{2}^{2}$.
@@
However, the application of these methods to generative models that use discrete diffusion models as priors is not straightforward. This limitation stems from two primary factors. First, with the former methods, diffusion models are assumed trained in the pixel domain. Second, while the latter methods can be extended to latent diffusion-type models, they encounter difficulties when handling discrete diffusion models, in which the generative process is inherently discrete. The core challenge lies in the lack of a direct mechanism to propagate gradients of the loss function through the generative process in discrete diffusion models. In such models, after generating discrete data, a non-differentiable operation (i.e., codebook assignment) is followed by a decoding operation into continuous space, which prevents the application of conventional gradient-based guidance.

### v2

# G2D2: Gradient-Guided Discrete Diffusion for Inverse Problem Solving

@@ Abstract
Recent literature has effectively leveraged diffusion models trained on continuous variables as priors for solving inverse problems. Notably, discrete diffusion models with discrete latent codes have shown strong performance, particularly in modalities suited for discrete compressed representations, such as image and motion generation. However, their discrete and non-differentiable nature has limited their application to inverse problems formulated in continuous spaces. This paper presents a novel method for addressing linear inverse problems by leveraging generative models based on discrete diffusion as priors. We overcome these limitations by approximating the true posterior distribution with a variational distribution constructed from categorical distributions and continuous relaxation techniques. Furthermore, we employ a star-shaped noise process to mitigate the drawbacks of traditional discrete diffusion models with absorbing states, demonstrating that our method performs comparably to continuous diffusion techniques with a lower GPU memory consumption. Our code is available at https://github.com/sony/g2d2.
@@ 1 Introduction
Diffusion models have gained significant attention as deep generative models, achieving remarkable success in image (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b; Dhariwal & Nichol, 2021; Esser et al., 2024), audio (Liu et al., 2023; Chen et al., 2024a), and video generation (Ho et al., 2022a; b). These models operate by iteratively corrupting data and then learning to reverse the corruption process, ultimately generating high-quality samples from noise. In parallel with continuous diffusion models, discrete diffusion models have emerged as a compelling alternative. These models have gained traction by demonstrating notable results not only in image (Gu et al., 2022), audio (Yang et al., 2023), and text generation (Austin et al., 2021; Lou et al., 2023a) but also in more specialized areas such as motion data (Lou et al., 2023b; Pinyoanuntapong et al., 2024), protein synthesis (Gruver et al., 2024), and graph generation (Vignac et al., 2023).
@@
Building on these advancements, researchers have made significant progress in expanding the application of diffusion models. They have explored using diffusion models, trained either directly in the pixel space or on latent representations derived from variational autoencoders (VAEs), to address inverse problems (Kawar et al., 2022; Chung et al., 2023b; Wang et al., 2023) and carry out various conditional generation tasks (Yu et al., 2023; Bansal et al., 2024; He et al., 2024) without the need for additional training. These efforts aim to use the powerful generative capabilities of diffusion models to tackle intricate problems and generate conditional outputs, all while preserving the models’ original trained parameters.
@@
The research on applying diffusion models to inverse problems and conditional generation has been primarily restricted to diffusion models trained in continuous spaces, and methods using pre-trained discrete diffusion models as priors remain limited (Gruver et al., 2024; Chen et al., 2024b; Li et al., 2024). One of the main reasons is that the inherent nature of the generation process in discrete diffusion models involves non-differentiable operations, posing a challenge for their application to inverse problems formulated in continuous spaces. Therefore, controlling discrete diffusion models often necessitates an additional trained network (Gruver et al., 2024; Nisonoff et al., 2024; Klarner et al., 2024; Vignac et al., 2023). Additionally, training-free methods have been confined to relatively low-dimensional data (Chen et al., 2024b) or to specific tasks such as image inpainting (Gu et al., 2022). Recent work by Singhal et al. (2025) proposes Feynman–Kac steering, a general inference-time framework for steering diffusion models with reward functions. While this approach does not require differentiating through reward functions and thus can be applied to discrete diffusion models, the reward functions they employ, such as human preference scores, impose weaker constraints on the generation process compared to typical inverse problems like deblurring, potentially significantly compromising consistency with measurements.
Despite these limitations in applying discrete diffusion models to inverse problems, their potential advantages in representing complex data distributions and generating high-fidelity samples motivate their exploration as priors. When examining existing approaches for continuous diffusion models, conventional approaches in continuous settings typically leverage gradient-based adjustments of the generation trajectory. These methods aim to refine intermediate latents by computing gradients of a likelihood loss function, ensuring they align well with the measurement equation or guidance target. This has been demonstrated in Chung et al. (2023b) and Yu et al. (2023). However, directly extending this gradient-based control to discrete diffusion models is challenging due to their inherently non-differentiable operations.
@@
Figure 1: Illustration of G2D2. At each time step $t$, variational categorical distribution $\tilde{p}_{\bm{\alpha}}$ is optimized with respect to the sum of prior loss and likelihood loss, followed by sampling $\mathbf{z}_{t-1}$. Both terms are continuously differentiable, enabling continuous optimization.
@@
To address the fundamental challenge of non-differentiability in discrete diffusion models, we propose Gradient-Guided Discrete Diffusion (G2D2), an inverse problem solving method that uses a discrete diffusion model as a prior. Our focus is on solving inverse problems using a generative model based on a discrete diffusion model, specifically designed for discrete latent variables such as those found in vector-quantized VAE (VQ-VAE) models. G2D2 bridges the gap between continuous and discrete domains by using a continuous relaxation technique to optimize the parameters of a variational distribution.
@@
In addition to the differentiability issue, discrete diffusion models present another challenge when used for inverse problems. These models often adopt “mask-absorbing” noise processes as their data corruption process due to generation quality. However, this mask-absorbing approach has a significant drawback in inverse-problem solving. For example, in discrete diffusion models for images, while a substantial portion of the image structure is determined in the initial stages of generation (i.e., when only a few tokens are determined), the mask-absorbing type does not allow transitions from an unmasked state to either a masked state or other unmasked states. Our experiments demonstrate that this restriction imposes a significant limitation on performance in solving inverse problems.
@@
To overcome this structural limitation of mask-absorbing processes, we incorporate the star-shaped noise process previously proposed in the context of continuous diffusion models (Okhotin et al., 2024; Zhang et al., 2024). This process removes the dependency between consecutive sampling steps, expanding the space that can be explored during generation. This process was originally proposed to enhance the performance of diffusion models (Okhotin et al., 2024), but was later introduced as a decoupled noise annealing process in the context of inverse problems using continuous diffusion models, demonstrating its effectiveness for continuous diffusion models (Zhang et al., 2024). In this study, we not only demonstrate that this process can be effectively applied to discrete diffusion models, but also discover that it uniquely addresses potential issues inherent in mask-absorbing-type discrete diffusion processes, specifically the inability to correct errors introduced in the early stages during later inference steps.
@@
To validate our proposed approach, we conduct comprehensive experiments comparing G2D2 to current methods using standard benchmark datasets. Our results demonstrate that G2D2 achieves comparable performance to continuous counterparts (within 0.020.020.050.05 LPIPS points) while reducing GPU memory usage by up to $77\%$ (4.74.7GiB vs 20.920.9GiB for PSLD). We also explore the application of a discrete prior-based motion-data-generation model to solve an inverse problem, specifically path-conditioned generation, without requiring further training. The results of our study indicate that G2D2 shows promise in tackling various inverse problems by leveraging pre-trained discrete diffusion models.
@@ 2.1 Discrete diffusion models for image generation
We first provide a brief overview of VQ-Diffusion (Gu et al., 2022; Tang et al., 2022), an image-generation model based on discrete diffusion processes. VQ-Diffusion generates images in a two-step process. It first produces discrete latent representations $\mathbf{z}_{0}$ using a discrete diffusion model trained on representations obtained from a pre-trained VQ-VAE model (van den Oord et al., 2017). It then transforms these representations into the continuous image space using a decoder. Each element of $\mathbf{z}_{0}\in\{1,\ldots,K\}^{d_{\mathbf{z}}}$ corresponds to one of the embedding vectors from the codebook, denoted as $\mathbf{B}:=\{\mathbf{b}_{1},\ldots,\mathbf{b}_{K}\},\mathbf{b}_{k}\in\mathbb{R}^{d_{\mathbf{b}}}$. During decoding, a variable $\mathbf{Z}\in\mathbf{B}^{d_{\mathbf{z}}}$ is constructed through codebook assignment, where $\left(\mathbf{Z}\right)_{i}=\mathbf{b}_{z_{0,i}}$ and $z_{0,i}$ denotes the $i$-th element of $\mathbf{z}_{0}$. This variable is then fed into a decoder $D:\mathbb{R}^{d_{\mathbf{b}}\times d_{\mathbf{z}}}\to\mathbb{R}^{d_{\mathbf{x}_{0}}}$ that maps from the discrete token embeddings to the continuous image space to obtain the final image: $\mathbf{x}_{0}=D(\mathbf{Z})$.
@@
where $\bm{v}(z_{t,i})\in\{0,1\}^{K+1}$ denotes a one-hot encoded vector representing the token at time step $t$, $K$ is the number of states from the VQ-VAE, and one additional state is for a special mask token [MASK]. These concepts will be formally introduced immediately below in the context of VQ-Diffusion. $Q_{t}\in\mathbb{R}^{(K+1)\times(K+1)}$ represents the transition matrix, which determines the probabilities of transitions between tokens. VQ-Diffusion uses a mask-absorbing-type forward process, which introduces a special masked token denoted as [MASK] in addition to the $K$ states from the VQ-VAE. The transition matrix is defined as follows, where the last column represents the transition probabilities for the [MASK] token:
@@
where the transition probabilities are determined by three parameters: $\alpha_{t}$, $\beta_{t}$, and $\gamma_{t}$. In this process, once a token transitions to the masked state, it remains masked in all subsequent steps, hence the term “absorbing.” Specifically, $\alpha_{t}$ represents the probability of a token remaining unchanged, $\beta_{t}$ denotes the probability of transitioning to a different unmasked token, and $\gamma_{t}$ indicates the probability of the token being replaced with the [MASK] token. These parameters satisfy the constraint $\alpha_{t}+(K-1)\beta_{t}+\gamma_{t}=1$ to ensure valid probability distributions. The probability $\beta_{t}$ between unmasked tokens is generally set to a very small value. These parameters are typically set so that $q(\mathbf{z}_{T}|\mathbf{z}_{0})$ assigns all probability mass to the [MASK] token, and we also adopt this assumption.
@@
where $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})$ represents the posterior distribution determined by the forward process, and $\tilde{p}_{\theta}$ denotes the denoising network that predicts the denoised token distribution at $t$. The output of $\tilde{p}_{\theta}$ is generally modeled as independent categorical distributions for each dimension in $\mathbf{z}_{0}$. In practical applications such as text-to-image generation, $\tilde{p}_{\theta}$ is trained with conditional information (e.g., text prompts in VQ-Diffusion), allowing for controlled generation. While the true data distribution $q(\mathbf{z}_{0})$ has dependencies across dimensions, the denoising network $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ typically models each dimension independently as categorical distributions. However, the complete reverse process defined in (3) implicitly captures some of these dependencies through the iterative application of the conditional distributions $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})$. We distinguish between two distributions: the clean distribution $\tilde{p}_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ directly estimated by the denoising network at a single step (which treats dimensions independently), and the distribution $p_{\theta}(\mathbf{z}_{0}|\mathbf{z}_{t})$ obtained by running the reverse diffusion process from $t$ to $t=0$, which better approximates the true data distribution with its dimensional dependencies.
@@ 2.2 Linear-inverse-problem settings
Inverse problems involve estimating unknown data from measurement. The relationship between the measurement $\mathbf{y}\in\mathbb{R}^{d_{\mathbf{y}}}$ and unknown ground-truth data $\mathbf{x}_{0}\in\mathbb{R}^{d_{\mathbf{x}_{0}}}$ can be represented as
@@
where $\mathbf{A}\in\mathbb{R}^{d_{\mathbf{y}}\times d_{\mathbf{x}_{0}}}$ is referred to as the forward linear operator, which describes the process by which the measurement $\mathbf{y}$ is obtained from data $\mathbf{x}_{0}$. We assume this operator is known. The term $\bm{\eta}$ represents measurement noise, which we assume follows an isotropic Gaussian distribution with a known variance $\sigma_{\bm{\eta}}^{2}$. Consequently, the likelihood function $q(\mathbf{y}|\mathbf{x}_{0})$ can be described as $\mathcal{N}(\mathbf{y};\mathbf{A}\mathbf{x}_{0},\sigma_{\bm{\eta}}^{2}\mathbf{I})$.
@@
One of the primary challenges in inverse problems is their ill-posed nature. This means that for any given measurement $\mathbf{y}$, multiple candidate solutions may exist. To address this issue and determine $\mathbf{x}_{0}$, a common approach is to assume a prior distribution for $\mathbf{x}_{0}$, such as a Laplace distribution. Diffusion models have been utilized as more powerful and expressive priors, offering enhanced capabilities in solving these inverse problems (Kawar et al., 2022; Chung et al., 2023b; Wang et al., 2023; Rout et al., 2023). These diffusion-based methods are able to produce data that not only fit the measurement data but also exhibit high likelihood under the prior model. Given a prior $q(\mathbf{x}_{0})$, the objective in the inverse problem is to sample from the posterior distribution $q(\mathbf{x}_{0}|\mathbf{y})$, which, according to Bayes’ theorem, is proportional to $q(\mathbf{y}|\mathbf{x}_{0})q(\mathbf{x}_{0})$.
@@
These methods can be categorized based on how they incorporate the information from the measurement $\mathbf{y}$ into the generation trajectory of diffusion models. Methods such as denoising diffusion restoration models (DDRM) (Kawar et al., 2022) and denoising diffusion null-space models (DDNM) (Wang et al., 2023) leverage the assumption of linear operators, using singular value decomposition of the forward process to control the generative process. In contrast, methods such as diffusion posterior sampling (DPS) (Chung et al., 2023b) and posterior sampling with latent diffusion (PSLD) (Rout et al., 2023) operate by propagating the gradient of a loss term through the generative process. This loss term is designed to maximize the measurement likelihood, specifically by minimizing the term $\|\mathbf{y}-\mathbf{A}\mathbf{x}_{0}\|_{2}^{2}$.
@@
However, while these methods work well with continuous diffusion models, their application to generative models that use discrete diffusion models as priors is not straightforward. This limitation stems from two primary factors. First, the former methods (DDRM and DDNM) are specifically designed for diffusion models trained directly in the pixel domain. Second, while the latter methods (DPS and PSLD) can be extended to latent diffusion models that operate in continuous latent spaces, they encounter difficulties when handling discrete diffusion models, where the generative process involves inherently non-differentiable operations. The core challenge lies in the lack of a direct mechanism to propagate gradients of the loss function through the generative process in discrete diffusion models. In such models, after generating discrete data, a non-differentiable operation (i.e., codebook assignment) is followed by a decoding operation into continuous space, which prevents the application of conventional gradient-based guidance.